# Argument-Aware Approach To Event Linking

### v1

@@ Abstract
Event linking connects event mentions in text with relevant nodes in a knowledge base (KB). Prior research in event linking has mainly borrowed methods from entity linking, overlooking the distinct features of events. Compared to the extensively explored entity linking task, events have more complex structures and can be more effectively distinguished by examining their associated arguments. Moreover, the information-rich nature of events leads to the scarcity of event KBs. This emphasizes the need for event linking models to identify and classify event mentions not in the KB as “out-of-KB,” an area that has received limited attention. In this work, we tackle these challenges by introducing an argument-aware approach. First, we improve event linking models by augmenting input text with tagged event argument information, facilitating the recognition of key information about event mentions. Subsequently, to help the model handle “out-of-KB” scenarios, we synthesize out-of-KB training examples from in-KB instances through controlled manipulation of event arguments. Our experiment across two test datasets showed significant enhancements in both in-KB and out-of-KB scenarios, with a notable 22% improvement in out-of-KB evaluations.
@@
I-Hung Hsuthanks:   The authors contribute ${}^{\;\;\ddagger}$    Zihan Xue11footnotemark: ${}^{\;\;\dagger}$   Nilay ${}^{\dagger}$    Sahil ${}^{\dagger}$ Premkumar ${}^{\ddagger}$    Jayanth ${}^{\mathsection}$    Nanyun ${}^{\dagger}$$^{\ddagger}$nformation Science Institute, University of Southern California ${}^{\dagger}$omputer Science Department, University of California, Los Angeles ${}^{\mathsection}$isco Research {ihunghsu, pnataraj}@isi.edu,   {zihanxue}@ucla.edu {jasriniv}@cisco.com,   {violetpeng}@cs.ucla.edu
@@ 1 Introduction
Event Linking Nothman et al. (2012); Ou et al. (2023) involves associating mentions of events in text with corresponding nodes in a knowledge base (KB). If no entry in the KB corresponds to the event mentioned, the model should predict it as out-of-KB, denoted as a “NIL” label, following prior works’ terminology Zhu et al. (2023); Dong et al. (2023); Yu et al. (2023a). This process of linking nodes can enhance text comprehension to facilitate various downstream applications, such as question-answering and recommendation systems Yu et al. (2023b); Li et al. (2020); Jacucci et al. (2021).
@@
We apply our design to a model architecture akin to Yu et al. (2023a); Pratapa et al. (2022) and conduct experiments on the two event linking datasets introduced by Yu et al. (2023a). Our approach yields a 22% accuracy improvement over previous baselines for out-of-KB testing and an over 1% increase for in-KB testing. Additionally, by comparing various methods for generating synthetic out-of-KB examples, we demonstrate that our data synthesis approach successfully balances in-KB and out-of-KB usage for event linking   [^1] We will release the code upon acceptance..
@@ 3 Method
3.1 Event Information Tagging
To make event linking models better capture event argument information, we use the UniST model  Huang et al. (2022) trained on the MAVEN dataset Wang et al. (2020) to first identify the event types of query events. Given predicted event types, we extract event arguments using the TagPrime model Hsu et al. (2023a); Huang et al. (2023) trained on the GENEVA dataset Parekh et al. (2023). Text with event argument tags is used by our model. Take the passage in Figure 3 as an example, “Germany” will be extracted as the “Assailant” of the invasion, “the Soviet Union” will be highlighted as the “Victim”. Overall, our system can support 115 event types with 220 different argument roles to cover the potential diverse type of events in the given queries. More relevant details about event trigger and argument extraction can be found in Appendix A.
@@ 3.2 Negative Data Generation
We address the challenge by leveraging our observation that events differ when argument configurations change. To generate a data point, we first sample an in-KB query from the training set, along with its tagged event mention and arguments. We then instruct a large language model (LLM) to adjust this example through a two-step process: first modifying the tagged event arguments and then making edits to ensure coherence and fluency, as demonstrated in Figure 3.  222 This two step generation is generated through a single prompt. We use GPT-3.5-Turbo (hhttps://platform.openai.com/docs/models/gpt-3-5-turbo) with 2-shot examples Brown et al. (2020) to instruct the model. More details about the prompt are listed in Appendix B.
To enhance the chance of generating more realistic out-of-KB query cases, we instruct the LLM to create context violating its own knowledge. However, there remains a possibility that the generated context may reference other events within the KB. To minimize this impact, in actual data usage, we treat our generated event query as a “negative” training data point when paired with top KB entries for the original sampled in-KB mention. Further details are provided in Section 3.3.
@@ 3.3 Model
After obtaining top-$k$ KB candidates $c_{1},c_{2},...,c_{k}$, a cross-encoder is employed to encode every pair $(q,c_{i})$ to a score $S(q,c_{i})$. The best candidate are selected by ranking the scores $c=\operatorname*{arg\,max}_{c_{i}}{S(q,c_{i})}$.
To handle out-of-KB scenarios, prior work Yu et al. (2023a), lacking out-of-KB training examples, generate final output $c_{\text{final}}$ by setting an arbitrary threshold $\theta$:
@@
Unlike this approach, our method introduces a learned “NIL” class trained with our synthetic negative data. During the re-ranking phase, we expand the candidate pool to include $k+1$ options, the extra one being a randomly initialized embedding that represents "NIL" :
@@ 4.1 Experimental Settings
  • The NYT dataset is a smaller, manually annotated test set. 2,500 lead paragraphs are sampled from The New York Times Annotated Corpus  [^3] https://catalog.ldc.upenn.edu/LDC2008T19 and then annotated through Amazon Mechanical Turk. The dataset comes from real-life news articles and contains out-of-KB event mentions that is not covered by Wikipedia.
@@
Evaluation Metrics We follow Yu et al. (2023a) to evaluate models using accuracy.
Baselines  [^4] We also explore LLM’s performance in Section 4.4.3.: (1) BM25, a term-based ranking algorithm for information retrieval. (2) BLINK Wu et al. (2020), the retrieve-and-rerank model architecture introduced for entity linking. We adopt its code but train it on our training set. The model takes pure query text as input. (3) EveLink Yu et al. (2023a), the current SOTA event linking model. It adopts BLINK but enhances the text query by adding local name entity information.
4.2 Implementation Details
During training, we first train the bi-encoder and retrieve the top 10 candidates using the trained bi-encoder. We then use the queries and their top 10 candidates to train the cross-encoder. During inference, we also first retrieve the top 10 candidates using the trained bi-encoder. Then, we re-rank the retrieved candidates using the trained cross-encoder and select the top 1 candidate to compare with the ground truth. To ensure fair comparison, all baselines and our methods follow this recipe. Results presented are the average of three random runs. Appendix C covers detailed implementation details.
@@ 4.2 Implementation Details
Table 2: Accuracy (%) on both Wikipedia (in-domain, in-KB) and NYT (out-of-domain, out-of-KB) test sets. The best performance is highlighted in bold. BLINK Wu et al. (2020); EveLink Yu et al. (2023a).
@@ 4.4 Analysis
In this section, we present studies to verify our two innovations, and explore other possibility of using LLMs on the task.
@@ 4.4.2 Effectiveness of Negative Data Generation
We benchmark our approach against two alternative methods for generating negative data to train the cross-encoder: (1) Non-argument-aware Data Generation, which also employs GPT-3.5-Turbo but does not incorporate event information into the prompts, as detailed in Appendix B; (2) KB Pruning, a strategy introduced by Dong et al. (2023) in the entity linking field. This method creates negative samples by randomly eliminating 10% of KB entries and marking the associated training data as negative examples external to the KB. Table 4 shows the comparison. We observe that while KB Pruning ensures high-quality negative examples outside the KB, it negatively affects performance on in-KB tests. In contrast, our method, designed with an emphasis on event information, effectively balances the use of in-KB and out-of-KB cases.
4.4.3 LLM Baseline Comparison
To the best of our knowledge, no existing event linking systems do not use LLMs yest. Thus, inspired by existing methods in information retrieval Pradeep et al. (2023), we explore the LLMs’ effectiveness by replacing the cross-encoder in the current pipeline with an instructed LLM to perform candidate re-ranking. We implement the baseline using GPT-3.5-Turbo and compare our proposed method with this LLM baseline.  [^5] Implementation details can be found in Appendix C We sample a subset of 1000 test examples and use the same bi-encoder system to ensure fair comparison.
From the comparison shown in Table 5, we observe that the method using using LLM to rerank is significantly inferior to our method. This underscores the advantage of our approach in utilizing LLM as an engine for generating negative data.
@@ 4.4.3 LLM Baseline Comparison
Table 3: Bi-encoder recall (%) on the Wikipedia test set. “R@1” stands for recall at 1, and so on. See Appendix D for more recall values.
Model Wiki. NYT Avg.
BLINK (no negative data usage) 78.74 27.13 52.94
  w/ Non-argument-aware method 79.09 54.08 66.59
  w/ KB Pruning 76.72 55.85 66.29
  w/ Argument-aware method (Ours) 79.22 55.18 67.20
Table 4: Analysis of alternative negative data generation methods. The best and the second-best are bolded and underlined, respectively
@@
5 Conclusion
Limitation
Broader Consideration
Appendix A Event Extraction Details
To accurately extract event arguments from the designated event in the input text, we first undertake event typing to categorize the input event into the most suitable event type. Subsequently, based on the assigned category, we identify and capture the event arguments along with their corresponding roles that are characteristic of that specific type.
For this purpose, we first employ UniST Huang et al. (2022) to conduct event typing. To accommodate the wide range of potential input event, we opt for MAVEN, a highly comprehensive event detection dataset in the general domain, comprising 168 event types. Subsequently, we train a UniST model on MAVEN to conduct event typing.
Upon identifying the event trigger and its corresponding type, we proceed with an event argument extraction model to obtain the related event arguments. We use TagPrime Hsu et al. (2023a), the current state-of-the-art in event argument extraction. To align with the MAVEN ontology, we adopt GENEVA, an extensive event argument extraction dataset, primarily based on the MAVEN ontology with minor modifications. These modifications result in a reduction of the final event types coverage to 115. Although other event extraction models exist, such as DEGREE Hsu et al. (2022), UniIE Lu et al. (2022), AMPERE Hsu et al. (2023b), and PAIE Ma et al. (2022), trained on various datasets, our approach represents the most comprehensive event extraction pipeline to date, encompassing the majority of event types. For additional information on event extraction models, we direct readers to the works of Peng et al. (2023); Huang et al. (2023) for a more thorough overview of available alternative methodologies.
Appendix B Data Generation Details
To generate negative data, we first get samples from the tagged training and validation data. We filter out the examples whose labeled event mentions are proper nouns or numeric to ensure the quality of our generated data. We also filter out the examples with less than two tagged event arguments to better apply our observation.
We then generate our negative data in a two-shot fashion with the prompt demonstrated in Figure 4.
@@ Appendix B Data Generation Details
Due to the cost of GPT-3.5-Turbo, we only generate 6600 examples generated from the training set and 1600 examples generated from the validation set in the end.
For the non-argument-aware data generation baseline, we employ the same method but without tagging event arguments. The prompt is adjusted to indicate the absence of event argument information. Figure 5 shows the actual prompt we use.
Note that the text content of the few shot examples is identical in the two cases. The only difference is the presence of event argument tags.
@@
Appendix C Implementation Details
We train all the baselines and our methods using Nvidia A100 40G GPUs. For biencoder training, it takes around 6 hrs to train. As for crossencoder training, the training time takes around 30 hrs.
For the bi-encoder, we use bert-base-uncased as our base encoder. We use a maximum query length of 300 and a maximum candidate length of 300. We train with a learning rate of 1e-5 and a batch size 48 for 15 epochs to best fit into our GPU usage.
For the cross-encoder, we also use bert-base-uncased as our base encoder. We use a maximum query length of 256 and a maximum candidate length of 256. We train with a learning rate of 2e-5 and batch size 6 for 20 epochs.
For the threshold $\theta$ used for baseline NIL prediction (see Section 3.3), we use follow the description in Yu et al. (2023a)’s paper and set it as $0.5$ after normalization.

### v2

@@ Abstract
Event linking connects event mentions in text with relevant nodes in a knowledge base (KB). Prior research in event linking has mainly borrowed methods from entity linking, overlooking the distinct features of events. Compared to the extensively explored entity linking task, events have more complex structures and can be more effectively distinguished by examining their associated arguments. Moreover, the information-rich nature of events leads to the scarcity of event KBs. This emphasizes the need for event linking models to identify and classify event mentions not in the KB as “out-of-KB,” an area that has received limited attention. In this work, we tackle these challenges by introducing an argument-aware approach. First, we improve event linking models by augmenting input text with tagged event argument information, which facilitates the recognition of key information about event mentions. Subsequently, to help the model handle “out-of-KB” scenarios, we synthesize out-of-KB training examples from in-KB instances through controlled manipulation of event arguments. Our experiments across two test datasets showed significant enhancements in both in-KB and out-of-KB scenarios, with a notable 22% improvement in out-of-KB evaluations.
@@
I-Hung Hsuthanks:   The authors contribute ${}^{\;\;\ddagger}$    Zihan Xue11footnotemark: ${}^{\;\;\dagger}$   Nilay Pochhi    Sahil Bansal Premkumar Natarajan    Jayanth Srinivasa§    Nanyun Peng Information Science Institute, University of Southern California Computer Science Department, University of California, Los Angeles §Cisco Research {ihunghsu, pnataraj}@isi.edu,   {zihanxue}@ucla.edu {jasriniv}@cisco.com,   {violetpeng}@cs.ucla.edu
@@ 1 Introduction
Event Linking Nothman et al. (2012); Ou et al. (2023) involves associating mentions of events in text with corresponding nodes in a knowledge base (KB). This process of linking nodes can enhance text comprehension to facilitate various downstream applications, such as question-answering and recommendation systems Yu et al. (2023b); Li et al. (2020); Jacucci et al. (2021).
@@
We apply our design to a model architecture akin to Yu et al. (2023a); Pratapa et al. (2022) and conduct experiments on the two event linking datasets introduced by Yu et al. (2023a). Our approach yields a 22% accuracy improvement over previous baselines for out-of-KB testing and an over 1% increase for in-KB testing. Additionally, by comparing various methods for generating synthetic out-of-KB examples, we demonstrate that our data synthesis approach successfully balances in-KB and out-of-KB usage for event linking. Our code can be found in https://github.com/PlusLabNLP/Argu_Event_Linking.
@@ 2 Related Work
3 Problem Statement
Event linking connects a mention of an event in the text to relevant entries within a knowledge base. If no matching entry exists, models predict “NIL,” as established in prior research Zhu et al. (2023); Yu et al. (2023a). The event-linking model analyzes a sequence of tokens $\mathbf{x}=x_{1},x_{2},...,x_{eve},...,x_{l}$, where $x_{eve}$ signifies the target event mention. Its objective is to output $y$, which can be either an entry $e_{i}$ from the knowledge base (size $n$), or “NIL” if no suitable match is found.
@@ 4 Method
4.1 Event Argument Tagging
To make event linking models better capture event argument information, we use the UniST model  Huang et al. (2022) trained on the MAVEN dataset Wang et al. (2020) to first identify the event types of query events. Given predicted event types, we extract event arguments using the TagPrime model Hsu et al. (2023a) trained on the GENEVA dataset Parekh et al. (2023). Text with event argument tags is used by our model. Take the passage in Figure 3 as an example, “Germany” will be extracted as the “Assailant” of the invasion, “the Soviet Union” will be highlighted as the “Victim”. More relevant details about event trigger and argument extraction can be found in Appendix B.
@@ 4.2 Negative Data Generation
We address the challenge by leveraging our observation that events differ when argument configurations change. To generate a data point, we first sample an in-KB query from the training set, along with its tagged event mention and arguments. We then instruct a large language model (LLM) to adjust this example through a two-step process: first modifying the tagged event arguments and then making edits to ensure coherence and fluency, as demonstrated in Figure 3.  111 This two step generation is generated through a single prompt. We use GPT-3.5-Turbo Ouyang et al. (2022) with 2-shot examples Brown et al. (2020) to instruct the model. More details about the prompt are listed in Appendix C.
To increase the likelihood of generating more realistic out-of-KB query cases, we instruct the LLM to create contexts that violates its own knowledge. However, there remains a possibility that the generated context may reference other events within the KB. To minimize this impact, in actual data usage, we treat our generated event query as a “negative” training data point when paired with top KB entries for the original sampled in-KB mention. Further details are provided in Section 4.3.
@@ 4.3 Model
After obtaining top-$k$ KB candidates $c_{1},c_{2},...,c_{k}\ \ \forall c_{i}\in\{e_{1},e_{2},...e_{n}\}$, a cross-encoder is employed to encode every pair $(q,c_{i})$ to a score $S(q,c_{i})$. The best candidate is selected by ranking the scores $c=\operatorname*{arg\,max}_{c_{i}}{S(q,c_{i})}$.
To handle out-of-KB scenarios, prior work Yu et al. (2023a), lacking out-of-KB training examples, generates the final output $c_{\text{final}}$ by setting an arbitrary threshold $\theta$:
@@
Unlike this approach, our method introduces a learned “NIL” class trained with our synthetic negative data. During the re-ranking phase, we expand the candidate pool to include $k+1$ options, the extra one being a randomly initialized embedding that represents “NIL” :
@@ 5.1 Experimental Settings
  • The NYT dataset is a smaller, manually annotated test set. 2,500 lead paragraphs are sampled from The New York Times Annotated Corpus and then annotated through Amazon Mechanical Turk. The dataset comes from real-life news articles and contains out-of-KB event mentions that is not covered by Wikipedia.
@@
Evaluation Metrics. We follow Yu et al. (2023a) to evaluate models using accuracy.
Compared Methods:
  • BM25: a term-based ranking algorithm for information retrieval.
  • GENRE: a generation-based entity linking model retrieves entities by generating their unique names. We follow the analysis from Yu et al. (2023a), who train GENRE on the event-linking training set and perform inference.
  • BLINK Wu et al. (2020): the retrieve-and-rerank model architecture introduced for entity linking. We adopt its code but train it on the event linking training set. BLINK processes the entire token sequence, enriched with special tokens to mark the beginning and end of the event mention, i.e., $x_{1},x_{2},...,[M_{s}],x_{eve},[M_{e}],...,x_{l}$, and apply retrieve-and-rerank training and inference. Here, $[M_{s}]$ and $[M_{e}]$ are the special markers.
  • EveLINK Yu et al. (2023a): the current SOTA event linking model that also follows retrieve-and-rerank framework. It adopts BLINK but enhances the text query by adding local name entity information. Specifically, EveLink initially performs named entity recognition on $\mathbf{x}$, identifying named entities, $ne_{1},ne_{2},...$ and their type $t_{1},t_{2},...$. It then combines the token and entity sequences as input to models, i.e., $x_{1},x_{2},...,[M_{s}],x_{eve},[M_{e}],...,x_{l},$ $[SEP],[t_{1s}],ne_{1},[t_{1e}],[t_{2s}],ne_{2},[t_{2e}],...$, where $[SEP]$ is the BERT separator, and $[t_{1s}]$, $[t_{1e}]$, … are the special tokens indicating the start and end of named entity types, respectively.
  • Our method also employs the retrieve-and-rerank framework. Yet, our method differs from BLINK in that we incorporate event argument information of $\mathbf{x}$ and apply negative data generation to augment the training data for better support on “out-of-KB” cases. Specifically, our model’s input format can be represented as $x_{1},x_{2},...,[r_{1s}],x_{arg1},[r_{1e}],...,[M_{s}],x_{eve},[M_{e}],$ $,[r_{2s}],x_{arg2},[r_{2e}],...$, where $x_{arg}$ denotes the event argument token, and $[r_{1s}]$, $[r_{1e}]$, …, are special tokens indicating the corresponding argument roles.
All reported results in this section are the average of three random runs. Appendix D covers all implementation details of the compared methods.
@@
Table 2: Accuracy (%) on both Wikipedia (in-domain, in-KB) and NYT (out-of-domain, out-of-KB) test sets. The best performance is highlighted in bold. We report GENRE Cao et al. (2021)’s numbers using the results from Yu et al. (2023a).
@@ 5.3 Analysis
In this section, we present studies to verify our two innovations. The exploration of the possibilities of using LLMs in event-linking modeling can be found in Appendix A.
@@ 5.3.2 Effectiveness of Negative Data Generation
We benchmark our approach against two alternative methods for generating negative data to train the cross-encoder: (1) Non-argument-aware Data Generation, which also employs GPT-3.5-Turbo but does not incorporate event information into the prompts, as detailed in Appendix C; (2) KB Pruning, a strategy introduced by Dong et al. (2023) in the entity linking field. This method creates negative samples by randomly eliminating 10% of KB entries and marking the associated training data as negative examples external to the KB. Table 4 shows the comparison. While KB Pruning ensures high-quality negative examples outside the KB, it negatively affects performance on in-KB tests. In contrast, our method, designed to emphasize event information, effectively balances the use of in-KB and out-of-KB cases.
@@
Table 3: Bi-encoder recall (%) on the Wikipedia test set. “R@1” stands for recall at 1, and so on. See Appendix E for more recall values.
Model Wiki. NYT
BLINK (no negative data usage) 78.74 27.13
w/ Non-argument-aware method 79.09 54.08
w/ KB Pruning 76.72 55.85
w/ Argument-aware method (Ours) 79.22 55.18
Table 4: Analysis of alternative negative data generation methods. The best and the second-best are bolded and underlined, respectively. Our method shows the best balance between in-KB and out-of-KB cases.
6 Conclusion
Acknowledgements
We gratefully acknowledge the insightful feedback provided by both the anonymous reviewers and the UCLA PLUS Lab and UCLA-NLP group members on earlier versions of this paper. This research was funded by a Cisco Sponsored Research Award.
Limitation
Broader Consideration
Appendix A Experiments with Large Language Models as Ranking Model
To the best of our knowledge, current event linking systems lack the integration of Large Language Models (LLMs). Inspired by recent information retrieval techniques leveraging LLMs Pradeep et al. (2023), we explore their effectiveness by incorporating an instructed LLM for candidate re-ranking within the existing event linking pipeline, replacing the traditional cross-encoder.
We establish a baseline using GPT-3.5-Turbo and compare our proposed method with this LLM baseline  [^2] Implementation details can be found in Appendix D.. To ensure a fair comparison, we utilize the same bi-encoder system and sample a subset of 1,000 test examples.
The results presented in Table 5 demonstrate that the method solely employing LLM for re-ranking significantly underperforms our approach. This finding highlights the key advantage of our method, which leverages LLMs for generating negative data, leading to superior performance.
@@ Appendix A Experiments with Large Language Models as Ranking Model
Appendix B Event Extraction Details
Our system employs a two-step approach for comprehensive event extraction from text. First, we perform event typing to categorize the event described in the input text. We leverage UniST Huang et al. (2022) to achieve this, trained on the MAVEN dataset, which offers broad coverage with 168 event types. This ensures our system can handle a wide range of potential input events.
Next, upon identifying the event type, we proceed with event argument extraction. We utilize TagPrime Hsu et al. (2023a), the state-of-the-art model in this domain. To maintain consistency with MAVEN’s ontology, we employ the GENEVA dataset, which aligns closely with MAVEN while encompassing 115 event types due to minor modifications. Consequently, our system offers comprehensive support for 115 event types, encompassing 220 distinct argument roles, effectively covering diverse event types within queries on our experimental dataset.
While alternative event extraction models exist, such as DEGREE Hsu et al. (2022), UniIE Lu et al. (2022), AMPERE Hsu et al. (2023b), and PAIE Ma et al. (2022), each trained on various datasets, our approach stands out as the most comprehensive event extraction pipeline to date. It captures the vast majority of event types, providing a robust solution for diverse user needs. For a more in-depth exploration of alternative methodologies, we recommend referring to the works of Peng et al. (2023); Huang et al. (2023).
Appendix C Data Generation Details
Our negative data generation begins by sampling from our tagged training and validation sets. To ensure the quality of the generated data, we filter out examples where the labeled event mentions are proper nouns or numeric values. Additionally, to better leverage our observations, we exclude examples with fewer than two tagged event arguments.
Following this filtering step, we employ a two-stage process to generate negative data, as detailed in the prompt provided in Figure 4.
@@ Appendix C Data Generation Details
Due to the high cost of GPT-3.5-Turbo, we only generated a final set of 6,600 examples from the training set and 1,600 examples from the validation set to train and develop our method.
We employed the same data generation method for the non-argument-aware baseline, but without tagging event arguments. The prompt was adjusted to reflect the absence of this information (see Figure 5 for the specific prompt used). Notably, the text content of the few-shot examples remained identical in both cases; the only distinction lies in the presence of event argument tags.
@@
Appendix D Implementation Details
Model Training and Inference Pipeline
Our system follows a two-stage training approach. First, we train a bi-encoder to encode queries and candidate event mentions. We then use the queries and their top 10 candidates to train the cross-encoder. During inference, we also first retrieve the top 10 candidates using the trained bi-encoder. Then, we re-rank the retrieved candidates using the trained cross-encoder and select the top 1 candidate to compare with the ground truth. To ensure a fair comparison, all baselines and our methods follow this recipe.
Training is conducted on Nvidia A100 40G GPUs. The bi-encoder typically requires approximately 6 hours to train, while the cross-encoder takes around 30 hours.
Model Architectures
Both the bi-encoder and cross-encoder leverage the pre-trained BERT-base-uncased model. For the bi-encoder, we set a maximum query and candidate length of 300 tokens. The training utilizes a learning rate of $51e-$ 5, a batch size of $48$, and runs for $15$ epochs to optimize GPU memory usage.
The cross-encoder also employs BERT-base-uncased, with a maximum query and candidate length of $256$ tokens. Here, the training process utilizes a learning rate of $52e-$ 5, a batch size of $6$, and runs for $20$ epochs.
Baseline Configurations
For the threshold $\theta$ used for baseline NIL prediction (see Section 4.3), we follow the description in Yu et al. (2023a)’s paper and set it as $0.5$ after normalization.